OpenStudy (anonymous):

a partical moves horizontally in a uniform circular motion ,over a horizontal xy plane .At one instant ,it moves through the point at coordintate (4m , 4m) with a velocity of -5.0i m/s and an acceleration of +12.5j m/s(sqr) . what is the (a)x and (b) y coordinates of the center of the circular path?

OpenStudy (anonymous):

@ghazi @vikrantg4

OpenStudy (anonymous):

horizontally and uniform circular ? o_O XD

OpenStudy (ghazi):

horizontally means on a plane

OpenStudy (ghazi):

done...

OpenStudy (ghazi):

\[r= \frac{ (v)^2 }{ a }\] = 5^2/12.5=2 m

OpenStudy (ghazi):

y coordinate of the center is 6m

OpenStudy (anonymous):

LOL OMG... how couldn't i have thought it

OpenStudy (ghazi):

x coordinate of the center is 4 straight up ...so find the radius and you'll have y coordinate as 2+4=6m

OpenStudy (ghazi):

@Yahoo! clear?

OpenStudy (anonymous):

thxxx...

OpenStudy (ghazi):

:)